# Logistic-Beta Processes for Dependent Random Probabilities with Beta Marginals

### v2

@@ 4.2 Logistic-beta dependent Dirichlet process mixture model
We model a covariate-dependent response density as $f(y\mid x)=\int\mathcal{K}(y\mid\theta)\mathrm{d}G_{x}(\theta)$, where $\{G_{x}\}$ follows a logistic-beta dependent Dirichlet process and $\mathcal{K}(\cdot\mid\theta)$ is a mixture component distribution parametrized by $\theta$. For example, we choose a normal kernel $\mathcal{K}(y\mid\theta)=N_{1}(y;\mu,\tau^{-1})$ and atom processes $\theta_{h}(x)=\{\mu_{h}(x),\tau_{h}(x)\}$ with $\mu_{h}(x)=\beta_{0h}+\beta_{1h}x$, $\tau_{h}(x)=\tau_{h}$,
$f(y_{i}\mid x_{i})$ $=\sum_{h=1}^{\infty}\Big{(}\sigma\{\eta_{h}(x_{i})\}\prod_{l<h}[1-\sigma\{\eta_{l}(x_{i})\}]\Big{)}N_{1}(y_{i};\mu_{h}(x_{i}),\tau_{h}^{-1})\quad(i=1,\dots,n),$ (12)
$\eta_{h}(\cdot)$ $\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}{\textsc{lbp}}(1,b,\mathcal{R}),\,\,\,\,(\beta_{0h},\beta_{1h})^{\mathrm{\scriptscriptstyle T}}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}N_{2}(0_{2},\Sigma_{\beta}),\quad\tau_{h}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Ga}(a_{\tau},b_{\tau})\,\,\,\,(h\in\mathds{N}),$ (13)
where logistic-beta processes in (13) and atom processes both induce dependence on $x$, leading to a mixture of normal linear regression models with covariate-dependent weights.
To adapt Algorithm 9 to model (12)-(13), we rely on the continuation-ratio logit representation of the multinomial distribution (Tutz,, 1991), which is also used in other stick-breaking process mixture models (Rodríguez and Dunson,, 2011; Rigon and Durante,, 2021). Denoting the mixture component allocations as $\{s_{i}\}_{i=1}^{n}$ and $\mathscr{Z}_{h}=\{z_{ih}:s_{i}>h-1\}$ as a set of binary variables such that $z_{ih}=1$ if $s_{i}=h$ and $z_{ih}=0$ if $s_{i}>h$, $\mathscr{Z}_{h}$ corresponds to observations of the latent logistic-beta process model for binary data with indices $I_{h}=\{i:s_{i}>h-1\}$, independently across $h=1,2,\dots$. We work on a truncated version having maximum number of components $H$ with $\sigma\{\eta_{H}(\cdot)\}\equiv 1$; if all the components are occupied we recommend increasing $H$.
@@
$\operatorname*{pr}(s_{i}=h\mid-)\propto\sigma\{\eta_{h}(x_{i})\}\prod_{l=1}^{h-1}[1-\sigma\{\eta_{l}(x_{i})\}]\mathcal{K}(y_{i}\mid\theta_{h}(x_{i})).$
[2] For $h=1$ to $h=H-1$, with index set $I_{h}=\{i:s_{i}>h-1\}$, $\lambda_{h}$ and $\eta_{h}(\cdot)$,
@@
[3] Sample component-specific parameters $(\beta_{0h},\beta_{1h},\tau_{h})^{\mathrm{\scriptscriptstyle T}}$ for $h=1,\dots,H$,
$\pi(\beta_{0h},\beta_{1h},\tau_{h}\mid-)\propto G^{0}(\beta_{0h},\beta_{1h},\tau_{h})\prod_{i:s_{i}=h}\mathcal{K}(y_{i}\mid\theta_{h}(x_{i})).$
Algorithm 2 One cycle of the blocked Gibbs sampler for the mixture model (12)–(13).
@@ 5 Empirical demonstrations
5.1 Simulation studies
Our simulation studies have two main objectives. Firstly, we aim to assess the benefits of posterior computation strategies involved in latent logistic-beta process models. We focus on how the blocking strategy in the Gibbs sampler and adaptive Pólya proposal affect mixing. Secondly, we compare the latent logistic-beta process with Gaussian copula models, which also give dependent random probabilities with beta marginals. We consider a nonparametric binary regression problem on a spatial domain $\mathscr{X}=[0,1]^{2}$, where spatial coordinates $\{x_{i}\}_{i=1}^{n}$ are generated uniformly at random from $\mathscr{X}$.
We considered the latent logistic-beta process model (8) with $(a,b)=(1,2)$ and the Gaussian copula model as two different data generation models. The Gaussian copula model first draws a sample $\big{\{}\zeta(x_{1}),\dots,\zeta(x_{n})\big{\}}^{\mathrm{\scriptscriptstyle T}}$ from a mean zero unit variance Gaussian process and then applies transformation $\zeta\mapsto F_{B}^{-1}\{F_{Z}(\zeta)\}$ to yield beta marginals, where $F_{Z}$ is the standard normal c.d.f. and $F_{B}^{-1}$ is the inverse c.d.f. of the Beta($1,2$) distribution. For both the latent logistic-beta process and Gaussian copula models, we used Matérn correlation kernels $\mathcal{R}_{M}(\cdot,\cdot;\varrho,\nu=1.5)$. We considered different degrees of spatial dependence $\varrho\in\{0.1,0.2,0.4\}$, resulting in six different data generation scenarios. Given the success probabilities, binary data $\{z(x_{i})\}_{i=1}^{n}$ of size $n=500$ were independently sampled, with $n_{\mathrm{test}}=100$ data points reserved to evaluate predictive performance. We simulated 100 replicated datasets for each data generation scenario.
@@ 5.1 Simulation studies
Table 1: Simulation results under various posterior inference algorithm settings. Entries are averages over 100 replicates, with Monte Carlo standard errors shown in parentheses. LBP, logistic-beta process; ESS, effective sample size of $\lambda$ based on 1000 posterior samples; ESS/sec, ESS divided by running time in seconds.
Data generation Algorithm settings ESS ESS/sec % Acceptance rate
Latent LBP, $\varrho=0.1$ Blocked Adapted 245.08 (12.86) 3.35 (0.18) 54.28 (1.03)
@@
Non-blocked Adapted 7.89 (0.35) 0.14 (0.01) 13.60 (0.26)
Non-adapted 7.13 (0.37) 0.13 (0.01) 12.44 (0.31)
Latent LBP, $\varrho=0.2$ Blocked Adapted 257.01 (16.32) 2.89 (0.18) 62.26 (1.12)
@@
Non-blocked Adapted 7.31 (0.32) 0.11 (0.00) 13.42 (0.22)
Non-adapted 6.58 (0.32) 0.10 (0.00) 12.48 (0.31)
Latent LBP, $\varrho=0.4$ Blocked Adapted 368.26 (17.54) 4.12 (0.20) 66.12 (0.98)
@@
Non-blocked Adapted 6.40 (0.29) 0.09 (0.00) 13.01 (0.22)
Non-adapted 6.56 (0.32) 0.10 (0.00) 12.88 (0.31)
Table 1 reports the performance of four different latent logistic-beta process posterior inference algorithm configurations based on the mixing behaviour of $\lambda$. “Non-blocked” refers to sampling $\lambda$ from $\pi(\lambda\mid\omega,\eta,z_{1:n})$ instead of $\pi(\lambda\mid\omega,z_{1:n})$, and “non-adapted” corresponds to the fixed independent Metropolis-Hastings proposal distribution of $\lambda$ using prior distribution $\mathrm{Polya}(a,b)$. We consider effective sample size based on 1,000 posterior samples, effective sample rate, and an acceptance rate of $\lambda$ from an independent Metropolis-Hastings sampler where a higher acceptance rate is desirable (Robert and Casella,, 2004, §7.6.1). The blocking scheme has a substantial impact on the mixing behaviour, leading to an effective sample size increase of up to 50 times. The adaptive Pólya proposal scheme improves the acceptance rate of $\lambda$ in all scenarios.
Table 2: Simulation results comparing latent logistic-beta process and Gaussian copula models. Entries are averaged over 100 replicates, with Monte Carlo standard errors in parentheses. LBP, logistic-beta process; RMSE, root mean squared error; CRPS, continuous ranked probability score; mESS/sec, multivariate effective sample size of latent success probabilities divided by running time in seconds.
@@
Gauss. copula, $\varrho=0.1$ Latent LBP 11.93 (0.14) 12.32 (0.17) 6.59 (0.10) 6.80 (0.11) 21.11 (0.21)
Gauss. copula 11.82 (0.13) 12.24 (0.16) 6.48 (0.09) 6.71 (0.11) 0.48 (0.01)
Gauss. copula, $\varrho=0.2$ Latent LBP 8.67 (0.15) 8.80 (0.16) 4.78 (0.10) 4.85 (0.10) 17.58 (0.16)
Gauss. copula 8.61 (0.16) 8.75 (0.17) 4.74 (0.10) 4.82 (0.11) 0.40 (0.01)
Gauss. copula, $\varrho=0.4$ Latent LBP 6.11 (0.16) 6.14 (0.16) 3.39 (0.10) 3.41 (0.10) 17.41 (0.17)
Gauss. copula 6.10 (0.16) 6.13 (0.16) 3.38 (0.10) 3.40 (0.10) 0.47 (0.01)
@@
Regarding sampling efficiency of latent success probabilities measured by multivariate effective sample size (Vats et al.,, 2019) divided by the algorithm’s running time in seconds, the latent logistic-beta process model dramatically outperforms the Gaussian copula model due to the significantly longer running time of the Gaussian copula model. Although the implementation of the Gaussian copula model with the Stan program is based on C++ code with highly optimized algorithms, the computation of the log-posterior and its gradient at each step demands high computational costs. The algorithm of the latent logistic-beta process model is currently implemented using R software, which has the potential for improvement in running time with a lower-level programming language. Appendix D contains further details on simulation settings and additional simulation results under different data generation scenarios.

### v3

@@ 4.3 Logistic-beta dependent Dirichlet process mixture model
We model a covariate-dependent response density as $f(y\mid x)=\int\mathcal{K}(y\mid\theta)\mathrm{d}G_{x}(\theta)$, where $\{G_{x}\}$ follows a LB-DDP and $\mathcal{K}(\cdot\mid\theta)$ is a mixture component distribution parametrized by $\theta$. For example, when we choose a normal kernel $\mathcal{K}(y\mid\theta)=N_{1}(y;\mu,\tau^{-1})$ and atom processes $\theta_{h}(x)=\{\mu_{h}(x),\tau_{h}(x)\}$, $\mu_{h}(x)=\beta_{0h}+\beta_{1h}x$, $\tau_{h}(x)=\tau_{h}$ with semi-conjugate base measure, the density regression model becomes
$f(y_{i}\mid x_{i})$ $=\sum_{h=1}^{\infty}\Big{(}\sigma\{\eta_{h}(x_{i})\}\prod_{l<h}[1-\sigma\{\eta_{l}(x_{i})\}]\Big{)}N_{1}(y_{i};\mu_{h}(x_{i}),\tau_{h}^{-1}),$ (14)
$\eta_{h}(x)$ $\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}{\textsc{lbp}}(1,b,\mathcal{R}),\,\,\,\,(\beta_{0h},\beta_{1h})^{\mathrm{\scriptscriptstyle T}}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}N_{2}(\bm{0}_{2},\bm{\Sigma}_{\beta}),\quad\tau_{h}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Ga}(a_{\tau},b_{\tau})$ (15)
for $i=1,\dots,n$ and $h\in\mathds{N}$, where logistic-beta processes in (15) and atom processes both induce dependence on $x$, leading to a mixture of normal linear regression models with covariate-dependent weights.
Algorithm 2 describes the blocked Gibbs sampler to estimate the parameters of the above model. To adapt Algorithm 9 to model (14)-(15), we rely on the continuation-ratio logit representation of the multinomial distribution (Tutz,, 1991), which is also used in other stick-breaking process mixture models (Rodríguez and Dunson,, 2011; Rigon and Durante,, 2021). Denoting the mixture component allocations as $\{s_{i}\}_{i=1}^{n}$ and $\mathscr{Z}_{h}=\{z_{ih}:s_{i}>h-1\}$ as a set of binary variables such that $z_{ih}=1$ if $s_{i}=h$ and $z_{ih}=0$ if $s_{i}>h$, $\mathscr{Z}_{h}$ corresponds to observations of the latent logistic-beta process model for binary data with indices $I_{h}=\{i:s_{i}>h-1\}$, independently across $h=1,2,\dots$. We work on a truncated version having maximum number of components $H$ with $\sigma\{\eta_{H}(\cdot)\}\equiv 1$; if all the components are occupied we recommend increasing $H$.
@@
$\operatorname*{pr}(s_{i}=h\mid-)\propto\sigma\{\eta_{h}(x_{i})\}\prod_{l=1}^{h-1}[1-\sigma\{\eta_{l}(x_{i})\}]N_{1}(y_{i};\beta_{0h}+\beta_{1h}x_{i},\tau_{h}^{-1}).$
[2] For $h=1,\dots,H-1$, with index set $I_{h}=\{i:s_{i}>h-1\}$, $\lambda_{h}$ and $\eta_{h}(\cdot)$,
@@
[3] For $h=1,\dots,H$:
  1. 1.
    Sample component-specific regression parameters $(\beta_{0h},\beta_{1h}$) from
    $(\beta_{0h},\beta_{1h}\mid-)\,\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}N_{2}\left(\tau_{h}\mathbf{V}_{h,\beta}\mathbf{X}_{h}^{\mathrm{\scriptscriptstyle T}}\mathbf{y}_{h},\mathbf{V}_{h,\beta}\right),$
    where $\mathbf{V}_{h,\beta}=(\tau_{h}\mathbf{X}_{h}^{\mathrm{\scriptscriptstyle T}}\mathbf{X}_{h}+\bm{\Sigma}_{\beta}^{-1})^{-1}$ and $\mathbf{X}_{h}$ and $\mathbf{y}_{h}$ are a matrix and a vector containing
the rows $(1,x_{i})$ and the scalars $y_{i}$ for which $s_{i}=h$, respectively.
  • 2.
    Sample component-specific precisions $\tau_{h}$ from
    $(\tau_{h}\mid-)\,\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{Ga}\left(a_{\tau}+\sum_{i=1}^{n}\mathbbm{1}(s_{i}=h)/2,b_{\tau}+\sum_{i:s_{i}=h}(y_{i}-\beta_{0h}-\beta_{1h}x_{i})^{2}/2\right).$
  • Algorithm 2 One cycle of the blocked Gibbs sampler for model (14)–(15).
    4.4 Extensions beyond dependent Dirichlet processes
    Logistic-beta processes can be further incorporated in dependent Bayesian nonparametric models beyond dependent Dirichlet processes. Consider two sequences $(a_{h})_{h=1}^{\infty}$, $(b_{h})_{h=1}^{\infty}$ such that $\sum_{h=1}^{\infty}\log(1+a_{h}/b_{h})$ diverges to positive infinity (Ishwaran and James,, 2001). Then, given the correlation kernel $\mathcal{R}$ and independent atom processes $\{\theta_{h}(x):x\in\mathscr{X}\}$, $h\in\mathds{N}$ with $G^{0}_{x}$ the base measure, one can consider the general class of logistic-beta dependent stick-breaking processes as
    $G_{x}(\cdot)=\sum_{h=1}^{\infty}\Big{(}\sigma\{\eta_{h}(x)\}\prod_{l<h}\left[1-\sigma\{\eta_{l}(x)\}\right]\Big{)}\delta_{\theta_{h}(x)}(\cdot),\quad\eta_{h}(x)\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}{\textsc{lbp}}(a_{h},b_{h},\mathcal{R}),$ (16)
    for $h\in\mathds{N}$, using independent but non-identical logistic-beta processes. For example, if $a_{h}=1-\varsigma$ and $b_{h}=b+h\varsigma$ for $\varsigma\in[0,1)$ and $b>-\varsigma$, this becomes a logistic-beta dependent Pitman-Yor process. That is, marginally for each $x$, $G_{x}$ follows a Pitman-Yor process with parameters $(\varsigma,b)$, and $\varsigma=0$ reduces to LB-DDP.
    When considering mixture models, processes as in equation (16) may be preferred over a DDP mixture for situations when data exhibit strong heterogeneity across $x$ in terms of the number of clusters; see Bassetti et al., (2014) for a detailed comparison. In such cases, the above dependent Pitman-Yor process construction based on logistic-beta provides computational benefits, since the inference algorithm is essentially the same as Algorithm 2. Finally, estimation for the hyperparameters $a_{h}$ and $b_{h}$ is each component can be suitably performed via particle marginal Metropolis-Hastings without direct evaluation of the Pólya density, as in the standard LB-DDP mixture model; see Section S.2.3 of the Appendix for a description.
@@ 5 Empirical demonstrations
    5.1 Nonparametric binary regression simulation study
    Our simulation studies on nonparametric binary regression have two main objectives. Firstly, we aim to assess the benefits of posterior computation strategies involved in latent logistic-beta process models. We focus on how the blocking strategy in the Gibbs sampler and adaptive Pólya proposal affect mixing. Secondly, we compare the latent logistic-beta process with Gaussian copula models, which also give dependent random probabilities with beta marginals. We consider a nonparametric binary regression problem on a spatial domain $\mathscr{X}=[0,1]^{2}$, where spatial coordinates $\{x_{i}\}_{i=1}^{n}$ are generated uniformly at random from $\mathscr{X}$.
    We considered the latent logistic-beta process model (8) with $(a,b)=(1,2)$ and the Gaussian copula model as two different data generation models. The Gaussian copula model first draws a sample $\big{\{}\zeta(x_{1}),\dots,\zeta(x_{n})\big{\}}^{\mathrm{\scriptscriptstyle T}}$ from a mean zero unit variance Gaussian process and then applies transformation $\zeta\mapsto F_{B}^{-1}\{F_{Z}(\zeta)\}$ to yield beta marginals, where $F_{Z}$ is the standard normal c.d.f. and $F_{B}^{-1}$ is the inverse c.d.f. of the Beta($1,2$) distribution. For both the latent logistic-beta process and Gaussian copula models, we used Matérn correlation kernels with fixed smoothness $\nu=1.5$. We considered different degrees of spatial dependence $\varrho\in\{0.1,0.2,0.4\}$, resulting in six different data generation scenarios. Given the success probabilities, binary data $\{z(x_{i})\}_{i=1}^{n}$ of size $n=500$ were independently sampled, with $n_{\mathrm{test}}=100$ data points reserved to evaluate predictive performance. We simulated 100 replicated datasets for each data generation scenario.
@@ 5.1 Nonparametric binary regression simulation study
    Table 1: Nonparametric binary regression simulation results comparing computational benefits of adaptive Pólya proposal scheme in the blocked Gibbs sampler. Entries shown are averages over 100 replicates, with Monte Carlo standard errors shown in parentheses. ESS, effective sample size of $\lambda$; ESS/sec, ESS divided by running time in seconds.
    Data generation Algorithm settings ESS ESS/sec Acceptance rate of $\lambda$ (%)
    LBP, $\varrho=0.1$ Adapted 245.08 (12.86) 3.35 (0.18) 54.28 (1.03)
@@
    LBP, $\varrho=0.2$ Adapted 257.01 (16.32) 2.89 (0.18) 62.26 (1.12)
@@
    LBP, $\varrho=0.4$ Adapted 368.26 (17.54) 4.12 (0.20) 66.12 (0.98)
@@
    Table 1 reports the benefits of the adaptive Pólya proposal scheme based on the mixing behavior of $\lambda$. Here “non-adapted” corresponds to the fixed independent Metropolis-Hastings proposal distribution of $\lambda$ using prior distribution $\mathrm{Polya}(a,b)$. We consider effective sample size (ESS) based on 1,000 posterior samples, effective sample rate (ESS/sec), and an acceptance rate of $\lambda$ from an independent Metropolis-Hastings sampler where a higher acceptance rate is desirable (Robert and Casella,, 2004, §7.6.1). The adaptive Pólya proposal scheme improves the acceptance rate of $\lambda$ in all scenarios. In Section S.3.1 of the Appendix, we also show that the blocking scheme in Steps 2 and 3 in Algorithm 9, which is made possible due to the normal mixture representation of the logistic-beta, has a substantial impact on the mixing behavior, leading to significant improvement on the effective sample size of $\lambda$.
    Table 2: Nonparametric binary regression simulation results comparing latent logistic-beta process (LBP) and Gaussian copula models (GauCop). Entries are averaged over 100 replicates, with Monte Carlo standard errors in parentheses. RMSE, root mean squared error; CRPS, continuous ranked probability score; mESS/sec, multivariate effective sample size of latent success probabilities divided by running time in seconds.
@@
    GauCop, $\varrho=0.1$ LBP 11.93 (0.14) 12.32 (0.17) 6.59 (0.10) 6.80 (0.11) 21.11 (0.21)
    GauCop 11.82 (0.13) 12.24 (0.16) 6.48 (0.09) 6.71 (0.11) 0.48 (0.01)
    GauCop, $\varrho=0.2$ LBP 8.67 (0.15) 8.80 (0.16) 4.78 (0.10) 4.85 (0.10) 17.58 (0.16)
    GauCop 8.61 (0.16) 8.75 (0.17) 4.74 (0.10) 4.82 (0.11) 0.40 (0.01)
    GauCop, $\varrho=0.4$ LBP 6.11 (0.16) 6.14 (0.16) 3.39 (0.10) 3.41 (0.10) 17.41 (0.17)
    GauCop 6.10 (0.16) 6.13 (0.16) 3.38 (0.10) 3.40 (0.10) 0.47 (0.01)
@@
    Regarding sampling efficiency of latent success probabilities measured by multivariate effective sample size (Vats et al.,, 2019) divided by the algorithm’s running time in seconds, the latent logistic-beta process model dramatically outperforms the Gaussian copula model due to the significantly longer running time of the Gaussian copula model. Although the implementation of the Gaussian copula model with the Stan program is based on C++ code with highly optimized algorithms, the computation of the log-posterior and its gradient at each step demands high computational costs. The algorithm of the latent logistic-beta process model is currently implemented using R software, which has the potential for improvement in running time with a lower-level programming language. Section S.3.1 of the Appendix contains further details on simulation settings and additional simulation results under different data generation scenarios.
    5.2 Bayesian density regression simulation study
    We now assess the density regression performance of the LB-DDP mixture model using simulated data. Specifically, we compare the LB-DDP with the linear DDP (De Iorio et al.,, 2009), which shares the same model as (14)-(15) but $\sigma\{\eta_{h}(x_{i})\}$ are replaced with $V_{h}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Beta}(1,b)$. This is a single-weights DDP where the only atoms depend on $x$. Hence, the difference between the two lies in the logistic-beta process prior alone, allowing us to evaluate the impact of its additional flexibility explicitly. We also compare LB-DDP with the logit stick-breaking process (LSBP, Ren et al.,, 2011). While it is not a DDP, the LSBP shares similar characteristics with LB-DDP by placing logistic transformed latent Gaussian processes in stick-breaking ratios, effectively acting as surrogate for a dependent Dirichlet process with covariate-dependent weights.
    We consider two different data generation scenarios from two previous studies, which we refer to as Scenario A (Dunson et al.,, 2007) and Scenario B (Wade and Inacio,, 2025), respectively. The true conditional densities are defined as
    $p_{A}(y\mid x_{i})$ $=\exp(-2x_{i})N_{1}(y;x_{i},0.1^{2})+\{1-\exp(-2x_{i})\}N_{1}(y;x_{i}^{4},0.2^{2}),$ (17)
    $p_{B}(y\mid x_{i})$ =$=\begin{cases}N_{1}(y;0,0.2^{2}),&-2\leq x_{i}\leq 2,\\ N_{1}(y;2x_{i}-4,0.05^{2}),&2<x_{i}\leq 5,\\ N_{1}(y;6,(x_{i}-5)^{2}/15+0.01),&5<x_{i}\leq 10,\\ \end{cases}$ (18)
    and we generated covariates from $x_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Unif}(0,1)$ for Scenario A and $x_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Unif}(-2,10)$ for Scenario B. With two sample sizes $n\in\{500,2000\}$, we simulated 100 replicated datasets for both scenarios.
    In each Scenario, we adopt the following modeling choices. For the correlation kernel $\mathcal{R}$ in LB-DDP, we choose a normalized feature map kernel based on a natural cubic spline basis with $6$ degrees of freedom $\{\phi_{k}(x)\}_{k=1}^{6}$, defined on $[0,1]$ for Scenario A and $[-2,10]$ for Scenario B. The same set of basis functions are used in LSBP, by letting $\sigma^{-1}\{V_{h}(x)\}=\eta_{h}(x)=\bm{\phi}(x)^{\mathrm{\scriptscriptstyle T}}\bm{\alpha}_{h}$ and $\bm{\alpha}_{h}\sim N_{q}(\bm{0}_{q},\sigma_{\alpha}^{2}\mathbf{I}_{q})$. This is in line with previous stick-breaking process literature (Rigon and Durante,, 2021; Horiguchi et al.,, 2024; Wade and Inacio,, 2025). For the hyperparameter choice, we consider the case when $b\in\{0.2,1,2\}$ for linear DDP and LB-DDP, and $\sigma_{\alpha}^{2}\in\{10^{2},\pi^{2}/3,0.2^{2}\}$ for LSBP. These three different hyperparameter choices result in similar tie probabilities (11) with $x=x^{\prime}$ (the gray dash in center panel of Figure 4), which are approximately $0.84,0.50$, and $0.34$, respectively. These tie probabilities do not depend on $x$ under normalized feature maps. Our indicator of performance are the regression error $E[\{\widehat{E(y\mid x)}-E(y\mid x)\}^{2}]^{1/2}$ and the density error $E\{\int|\widehat{p(y\mid x)}-p(y\mid x)|\mathrm{d}y\}$ averaged across all test covariates $x$, similar to Wade and Inacio, (2025). See Section S.3.2 of the Appendix for further details.
    Figure 5: Density regression simulation results under two different data generation scenarios with sample size $n=500$ and three different hyperparameter settings based on prior tie probability. Density error and regression error quantify the accuracy of conditional density and conditional mean estimates averaged across test covariates, respectively.
    Figure 5 reports the result when $n=500$; boxplots for $n=2000$ are available in Figure S.3.1 of the Appendix. Compared to the linear DDP mixture model, the performance of the LB-DDP and LSBP mixture model are significantly improved in both scenarios, which shows a clear advantage of the flexibility offered by covariate-dependent weights. Meanwhile, compared to LB-DDP, the results of the LSBP mixture model are highly sensitive to the choice of hyperparameters, even though they are calibrated to induce similar prior tie probabilities. Thus, consistent with findings of Wade and Inacio, (2025), the choice of hyperparameters in stick-breaking process priors can drastically affect performance, partly due to an unknown covariate-dependent marginal structure of prior that can abruptly change in certain regions of $x$. Overall, the density regression simulation result shows the clear benefits of LB-DDP in terms of flexibility, interpretability, and robustness to the choice of hyperparameters.